# IMSRG with flowing 3 body operators, and approximations thereof

### v1

@@ II.2 Magnus formulation
The first term in (14) will scale as $A^{3}$. The second term will scale as $A^{2}N_{q}$, where $N_{q}$ is the number of quasiparticles  [^2] Here, we assume that the size of $\langle\Psi|\{a^{\dagger}a\}|\Psi\rangle$ is reasonably approximated by the diagonal terms $N_{q}=\sum_{p}\langle p^{\dagger}p\rangle+\sum_{h}\langle hh^{\dagger}\rangle$. In nuclear structure applications, many of the off-diagonal terms will be suppressed by selection rules on the quantum numbers $j,\pi,t_{z}$.. The third term will go as $AN_{q}^{2}$, and the fourth as $N_{q}^{3}$, so we see that we have an organization in powers of $N_{q}/A$ [30]. The approximation is improved by the IMSRG flow, which suppresses excitations so that $N_{q}(s)\to 0$ as $s\to\infty$. This argument is complicated somewhat by the fact that the three-body operators induced by the IMSRG flow are not uniformly distributed over all orbits, but instead are concentrated near the Fermi surface. Nevertheless, so long as the typical number of quasiparticles is smaller than the number of particles feeling the induced forces, we expect the effect of many-body operators to be suppressed.
& $[3,3]\to 0$ $[3,3]\to 1$ $[3,2]\to 1$ $[3,1]\to 2$ $[3,2]\to 2$ $[3,3]\to 2$ $[2,2]\to 3$ $[3,1]\to 3$ $[3,2]\to 3$ $[3,3]\to 3$
@@ III Approximations to full IMSRG(3)
The full commutator expressions necessary for IMSRG(3) in the uncoupled and $J$-coupled representations are presented in Appendices A and B respectively (see also refs [31, 32, 17, 33]). The corresponding Hugenholtz skeleton diagrams which enter at the IMSRG(3) level are presented in Fig. II.3. The naive computational scaling to evaluate a particular term can be read off of the diagram by counting the number of fermion lines. For example, the first diagram in Fig. II.3, labeled $[3,3]\to 0$, has six fermion lines and thus scales as $N^{6}$, where $N$ is the number of single-particle states in the basis. The final diagram in Fig. II.3 involves nine fermion lines and thus scales as $N^{9}$. This is too expensive for realistic applications (see section V), and so we shall explore some approximation schemes which can render the calculation more tractable.
@@ III.1 Goose-tank diagrams and the IMSRG(2) approximation
(a)     (b)           (c)           (d)
Figure 3: (a) and (b): Diagrammatic illustration of a contributions to $[\eta,[\eta,H]_{3}]_{2}$, which we denote goose-tank diagrams. (c) Diagram arising from $[\eta,[\eta,H]_{1}]_{2}$ which gives a contribution identical to that of (a). (d) Diagram arising from $[\eta,[\eta,H]_{1}]_{2}$ which is approximately equal to (b). In these diagrams, vertices of $H$ are represented by circles, while vertices of $\eta$ are represented by squares.
@@ III.2 Perturbative triples
that is, we should take half the value of the $[3,3]_{0}$ commutator [^4] Here we introduce the shorthand notation $[a,b]_{c}$ for the $c$-body piece of the commutator of an $a$-body operator with a $b$-body operator.. Taking the explicit form of the commutator, we have the perturbative correction to the energy [32]
@@ III.3 Truncation to $N^{7}$ scaling terms
Alternatively, we may assume that not all topologies contributing to the full IMSRG(3) commutator expression are equally important. A pragmatic way to approximate the full expression is to retain only those terms which scale as $N^{7}$ or better. Explicitly, this amounts to neglecting the following topologies shown in Fig. II.3: $[3,3]_{2}$, $[3,2]_{3}$, and $[3,3]_{3}$. In this approximation, errors made in the zero-body piece of $H(s)$ will be at least fifth order in the potential, errors to the one-body and two-body parts will be at least fourth order, and the error in the three-body part will be at least third order. This truncation scheme was previously explored in Refs. [39, 33], and we will refer to it as IMSRG(3N7). In principle we can also consider terms scaling as $N^{8}$ or better, but in practice we find these extra terms are not worth the substantial computational effort required for direct evaluation.
@@ IV.2 Lipkin-Meshkov-Glick model
Moving on to the IMSRG solution, we note that the LGM model was previously considered in a flow equation approach by Pirner and Friman [43], who used the quasi-spin operators (without normal ordering) to formulate the flow equation. They found it necessary to include an induced three-body term, but even then obtained results which were less accurate than second order perturbation theory. More recently, the flow equation approach was applied to obtain $1/N$ corrections to the thermodynamic limit [44]. Here we are interested in finite system, and in understanding the impact of truncations in the particle-rank of the flowing operators. Making an approximation equivalent to IMSRG(2) requires the use of normal-ordered operators. Commutators of normal-ordered strings of quasi-spin operators are presented in Appendix C. While we may construct the IMSRG flow equations in terms of the normal-ordered quasi-spin operators, integrating these equations must still be done numerically. However, in the Magnus formulation, we may directly solve for the Magnus operator at the fixed point, obtaining (see Appendix C)
@@
Figure 9: IMSRG flow in the LGM model, for the IMSRG(2) and IMSRG(2) approximations for several values of $V(0)$. Arrows indicate the direction of the flow. The red lines indicate the exact solutions from (34).
@@ Acknowledgements.
S.R.S. would like to thank Matthias Heinz for help on optimizing the implementation and finding a typo in the commutator expressions. B.C.H. was supported in part by the U.S. Department of Energy (DoE), Office of Science, under SciDAC-5 (NUCLEI collaboration) under grant DE-FG02-97ER41014. This research was supported in part by the Notre Dame’s Center for Research Computing.
@@ Appendix A Uncoupled commutator expressions
$+\sum_{a}P_{ij/k}P_{l/mn}(A_{ijla}B_{akmn}-B_{ijla}A_{akmn})$

### v2

@@ II.2 Magnus formulation
The first term in (14) will scale as $A^{3}$. The second term will scale as $A^{2}N_{q}$, where $N_{q}$ is the number of quasiparticles  [^2] Here, we assume that the size of $\langle\Psi|\{a^{\dagger}a\}|\Psi\rangle$ is reasonably approximated by the diagonal terms $N_{q}=\sum_{p}\langle p^{\dagger}p\rangle+\sum_{h}\langle hh^{\dagger}\rangle$. In nuclear structure applications, many of the off-diagonal terms will be suppressed by selection rules on the quantum numbers $j,\pi,t_{z}$.. The third term will go as $AN_{q}^{2}$, and the fourth as $N_{q}^{3}$, so we see that we have an organization in powers of $N_{q}/A$ [31, 32]. The approximation is improved by the IMSRG flow, which suppresses excitations so that $N_{q}(s)\to 0$ as $s\to\infty$. This argument is complicated somewhat by the fact that the three-body operators induced by the IMSRG flow are not uniformly distributed over all orbits, but instead are concentrated near the Fermi surface. Nevertheless, so long as the typical number of quasiparticles is smaller than the number of particles feeling the induced forces, we expect the effect of many-body operators to be suppressed.
@@ III Approximations to full IMSRG(3)
The full commutator expressions necessary for IMSRG(3) in the uncoupled and $J$-coupled representations are presented in Appendices A and B respectively (see also refs [33, 34, 17, 35]). The corresponding Hugenholtz skeleton diagrams which enter at the IMSRG(3) level are presented in Fig. 1. The naive computational scaling to evaluate a particular term can be read off of the diagram by counting the number of fermion lines. For example, the first diagram in Fig. 1, labeled [^3] Here we use the shorthand notation $[a,b]_{c}$ for the $c$-body piece of the commutator of an $a$-body operator with a $b$-body operator. $[3,3]_{0}$, has six fermion lines and thus scales as $N^{6}$, where $N$ is the number of single-particle states in the basis. The final diagram in Fig. 1 involves nine fermion lines and thus scales as $N^{9}$. This is too expensive for realistic applications (see section V), and so we shall explore some approximation schemes which can render the calculation more tractable.
@@ III.1 Goose-tank diagrams and the IMSRG(2) approximation
Figure 3: (a) and (b): Diagrammatic illustration of a contributions to $[\eta,[\eta,H]_{3}]_{2}$, which we denote “goose-tank” diagrams. (c) Diagram arising from $[\eta,[\eta,H]_{1}]_{2}$ which gives a contribution identical to that of (a). (d) Diagram arising from $[\eta,[\eta,H]_{1}]_{2}$ which is approximately equal to (b). In these diagrams, vertices of $H$ are represented by circles, while vertices of $\eta$ are represented by squares.
@@ III.2 Perturbative triples
that is, we should take half the value of the $[3,3]_{0}$ commutator. Taking the explicit form of the commutator, we have the perturbative correction to the energy [34]
@@ III.3 Truncation to $N^{7}$ scaling terms
Alternatively, we may assume that not all topologies contributing to the full IMSRG(3) commutator expression are equally important. A pragmatic way to approximate the full expression is to retain only those terms which scale as $N^{7}$ or better. Explicitly, this amounts to neglecting the following topologies shown in Fig. 1: $[3,3]_{2}$, $[3,2]_{3}$, and $[3,3]_{3}$. In this approximation, errors made in the zero-body piece of $H(s)$ will be at least fifth order in the potential, errors to the one-body and two-body parts will be at least fourth order, and the error in the three-body part will be at least third order. This truncation scheme was previously explored in Refs. [41, 35], and we will refer to it as IMSRG(3N7). In principle we can also consider terms scaling as $N^{8}$ or better, but in practice we find these extra terms are not worth the substantial computational effort required for direct evaluation.
@@ IV.2 Lipkin-Meshkov-Glick model
Moving on to the IMSRG solution, we note that the LGM model has been frequently studied within flow equation approach [45, 46, 31, 47, 48, 49], notably by Pirner and Friman [45], who used the quasi-spin operators (without normal ordering) to formulate the flow equation. They found it necessary to include an induced three-body term, but even then obtained results which were less accurate than second order perturbation theory. More recently, the flow equation approach was applied to obtain $1/N$ corrections to the thermodynamic limit [49]. Here we are interested in finite system, and in understanding the impact of truncations in the particle-rank of the flowing operators. Making an approximation equivalent to IMSRG(2) requires the use of normal-ordered operators. Commutators of normal-ordered strings of quasi-spin operators are presented in Appendix C. While we may construct the IMSRG flow equations in terms of the normal-ordered quasi-spin operators, integrating these equations must still be done numerically. However, in the Magnus formulation, we may directly solve for the Magnus operator at the fixed point, obtaining (see Appendix C)
@@
Figure 9: IMSRG flow in the LGM model, for the IMSRG(2) and IMSRG(2) approximations for several values of $V(0)$. Arrows indicate the direction of the flow. The red lines indicate the exact solutions from (34). cf Ref. [48].
@@ Acknowledgements.
S.R.S. would like to thank Matthias Heinz for help on optimizing the implementation and finding a typo in the commutator expressions. S.R.S. and B.C.H. acknowledge support from the U.S. Department of Energy (DoE), Office of Science, under SciDAC-5 (NUCLEI collaboration) under grant DE-FG02-97ER41014. This research was supported in part by the Notre Dame’s Center for Research Computing. This work benefitted from the use of the amc code [52] to validate the $J$-coupled expressions. The imsrg++ code [53] employed in this work uses the Armadillo library [54, 55].
@@ Appendix A Uncoupled commutator expressions
$-\sum_{a}P_{ij/k}P_{lm/n}(A_{ijna}B_{kalm}-B_{ijna}A_{kalm})$